OpenStudy (anonymous):

three point charges each 4 microcoulombs are placed at the three corners of a square of side 20cm. find the magnitude of the force on each.

OpenStudy (anonymous):

Hi welcome to OPenstudy! First thing you have to do is to draw a diagram:  Find R by using pythagorean theorem which is \(c^2=a^2+b^2\) After you solve for the distance R, calculate the \(F_E\) on each charge by using the formula: \(\Large F_E=\frac{kq_1q_2}{r^2}\). But the problem is you don't know if these charges are positive or negative.

OpenStudy (goformit100):

Hello, and A Warm Welcome to Open Study!